### v6

# Robust Tail Index Estimation under Random Censoring via Minimum Density Power Divergence

Abstract
We propose a robust estimator for the tail index of Pareto-type distributions under random right-censoring, constructed within the minimum density power divergence (MDPD) framework and based on the Nelson–Aalen estimator of the cumulative hazard function. To our knowledge, this is the first application of the MDPD methodology to tail index estimation in the presence of random censoring. Under mild regularity conditions and within the weak censoring regime, the estimator is shown to be consistent and asymptotically normal. Its finite-sample performance is assessed through Monte Carlo simulations, revealing improved robustness–efficiency trade-offs compared to standard non-robust tail index estimators. Robustness is investigated under both pre-censoring and post-censoring contamination schemes. While pre-censoring contamination provides a meaningful framework for robustness assessment, post-censoring contamination directly alters the observable data and highlights the sensitivity of reconstruction-based approaches. The practical relevance of the method is illustrated using an insurance claims dataset with light censoring and fully observable extremes. An additional application to AIDS survival data is included for illustrative purposes, emphasizing the challenges encountered under stronger censoring.
Keywords: Asymptotic normality; Heavy-tails; Robust estimation; Random right-censoring; Tail index.
AMS 2020 Subject Classification: 62G32; 62G05; 62G20; 62G35.
[…]
E-mail address:
nourelhouda.guesmia@univ-biskra.dz (N. Guesmia)
djamel.meraghni@univ-biskra.dz (D. Meraghni)
1. Introduction
Right-censored data pose a common challenge in statistics, as values are only known up to a certain threshold, while those beyond it remain unobserved. This situation frequently arises in fields such as survival analysis, reliability engineering, and insurance. Pareto-type distributions are commonly used to model data exhibiting extreme values. These distributions are characterized by a relatively high probability of producing observations far from the mean, in contrast with the normal distribution. They are particularly relevant in applications such as insurance, finance, and environmental studies, where rare events can have a substantial impact.
Let $X_{1},X_{2},...,X_{n}$ be a sample of size $n\geq 1$ from a random variable (rv) $X,$ and let $C_{1},C_{2},...,C_{n}$ be another sample from a rv $C,$ both defined on a probability space $\left(\Omega,\mathcal{A},\mathbf{P}\right),$ with continuous cumulative distribution functions (cdfs) $F$ and $G,$ respectively. Assume that $X$ and $C$ are independent. We also suppose that $X$ is right censored by $C,$ meaning that for each index $1\leq j\leq n,$ we can only observe the variable
$Z_{j}:=\min\left\{X_{j},C_{j}\right\}$
and the indicator variable
$\delta_{j}:=\mathbb{I}_{\left\{X_{j}\leq C_{j}\right\}},$
which determines whether or not $X$ has been observed. We assume that the tail functions $\overline{F}:=1-F$ and $\overline{G}:=1-G$ are regularly varying at infinity (or Pareto-type) with positive tail indices $\gamma_{1}>0$ and $\gamma_{2}>0,$ allowing for slowly varying deviations from the strict Pareto form. More precisely, for any $x>0,$
$\lim_{u\rightarrow\infty}\frac{\overline{F}(ux)}{\overline{F}(u)}=x^{-1/\gamma_{1}}\text{ and }\lim_{u\rightarrow\infty}\frac{\overline{G}(ux)}{\overline{G}(u)}=x^{-1/\gamma_{2}}.$ (1.1)
Equivalently, these tail behaviors can be expressed in terms of slowly varying functions $\ell_{1}\left(x\right)$ and $\ell_{2}\left(x\right)$ to account for deviations from strict Pareto distributions:
$\overline{F}(x)=x^{-1/\gamma_{1}}\ell_{1}\left(x\right)\text{ and }\overline{G}(x)=x^{-1/\gamma_{2}}\ell_{2}\left(x\right),$ (1.2)
where $\ell_{1}\left(x\right)$ and $\ell_{2}\left(x\right)$ vary slowly at infinity, i.e., for any fixed $t>0,$
$\ell_{j}\left(tx\right)/\ell_{j}\left(x\right)\rightarrow 1,\text{ as }x\rightarrow\infty,\text{ }j=1,2.$
These expressions make explicit the Pareto-type form of the distributions and clarify that the limit relations correspond to first-order asymptotics, while the slowly varying functions explicitly capture deviations from a strict Pareto model.
We denote the cdf of $Z$ by $H.$ Then, using the independence of $X$ and $C,$ we have
$\overline{H}=\overline{F}\times\overline{G},$
which implies that $\overline{H}$ is also regularly varying at infinity, with tail index
$\gamma:=\frac{\gamma_{1}\gamma_{2}}{\gamma_{1}+\gamma_{2}}.$
This characterization of the tail behavior of the observed variable $Z$ provides a firm basis for estimating the tail index $\gamma$ under right censoring. In the subsequent sections, we develop estimation procedures that appropriately account for censored observations and examine their asymptotic properties, including robustness to contamination.
In the presence of extreme values and right-censored data, various methods have been proposed for estimating the tail index. Several techniques have been developed to address the specific challenges posed by such data. Many studies have focused on modifying traditional tail index estimation methods, notably Hill’s estimator (Hill, 1975) to accommodate censored observations. In particular, Einmahl et al. (2008) adapted this estimator to handle right-censored data, leading to the following form:
$\widehat{\gamma}_{1}^{\left(\mathrm{EFG}\right)}:=\frac{\widehat{\gamma}^{\left(H\right)}}{\widehat{p}}$,
where
$\widehat{\gamma}^{\left(\mathrm{H}\right)}:=k^{-1}{\displaystyle\sum\limits_{i=1}^{k}}\log\left(Z_{n-i+1:n}/Z_{n-k:n}\right),$
denotes the classical Hill estimator corresponding to the tail index $\gamma,$ and
$\widehat{p}:=k^{-1}{\displaystyle\sum_{i=1}^{k}}\delta_{\left[n-i+1:n\right]},$
represents an estimator of the proportion of upper non-censored observations,
$p:=\frac{\gamma_{2}}{\gamma_{1}+\gamma_{2}}.$ (1.3)
The integer sequence $k=k_{n}$ represents the number of top order statistics used in the estimation of the tail index, such that $k\rightarrow\infty$ and $k/n\rightarrow 0$ as $n\rightarrow\infty.$ This ensures that enough extreme observations are included for reliable estimation, while avoiding the inclusion of too many moderate values that could bias the tail index estimation.
The sequence of rvs $Z_{1:n}\leq...\leq Z_{n:n}$ represents the order statistics pertaining to the sample $Z_{1},...,Z_{n}$ and $\delta_{\left[1:n\right]},...,\delta_{\left[n:n\right]}$ denotes the corresponding concomitant values, satisfying $\delta_{\left[j:n\right]}=\delta_{i}$ for $i$ such that $Z_{j:n}=Z_{i}.$ This notation clarifies the alignment between the ordered observations and their censoring indicators, which is essential for constructing the censored Hill-type estimator.
On the basis of a Kaplan-Meier integration, Worms and Worms (2014) proposed a consistent estimator to the tail index $\gamma_{1}$ defined by
$\widehat{\gamma}_{1}^{\left(\mathrm{W}\right)}:={\displaystyle\sum\limits_{i=1}^{k}}\frac{\overline{F}_{n}^{\left(KM\right)}\left(Z_{n-i:n}\right)}{\overline{F}_{n}^{\left(KM\right)}\left(Z_{n-k:n}\right)}\log\frac{Z_{n-i+1:n}}{Z_{n-i:n}},$
where
$F_{n}^{\left(\mathrm{KM}\right)}\left(x\right):=\underset{Z_{i:n}\leq x}{1-{\displaystyle\prod}}\left(\dfrac{n-i}{n-i+1}\right)^{\delta_{\left[i:n\right]}},$
denotes the popular Kaplan-Meier estimator of the cdf $F$ (Kaplan and Meier, 1958). The asymptotic normality of $\widehat{\gamma}_{1}^{\left(\mathrm{W}\right)}$ is established in Beirlant et al. (2019) by considering Hall’s model (Hall, 1982). Bias reduction in tail index estimation under censoring has been addressed in Beirlant et al. (2016), Beirlant et al. (2018) and Beirlant et al. (2019).
By using a Nelson-Aalen integration, more recently Meraghni et al. (2025) derived a new estimator for $\gamma_{1}$ given by
$\widehat{\gamma}_{1}^{\left(\mathrm{MNS}\right)}:={\displaystyle\sum\limits_{i=1}^{k}}\frac{\delta_{\left[n-i+1:n\right]}}{i}\frac{\overline{F}_{n}^{\left(NA\right)}\left(Z_{n-i+1:n}\right)}{\overline{F}_{n}^{\left(NA\right)}\left(Z_{n-k:n}\right)}\log\frac{Z_{n-i+1:n}}{Z_{n-k:n}},$
where
$F_{n}^{\left(\mathrm{NA}\right)}\left(x\right)=1-\prod_{Z_{i:n}<x}\exp\left\{-\frac{\delta_{\left[i:n\right]}}{n-i+1}\right\},$ (1.4)
denotes the well-known Nelson-Aalen estimator of the cdf $F$ (Nelson, 1972). Note that
$\frac{\overline{F}_{n}^{\left(\mathrm{NA}\right)}\left(Z_{n-i+1:n}\right)}{\overline{F}_{n}^{\left(\mathrm{NA}\right)}\left(Z_{n-k:n}\right)}=\prod_{j=i+1}^{k}\exp\left\{-\frac{\delta_{\left[n-j+1:n\right]}}{j}\right\},$ (1.5)
thus the formula of $\widehat{\gamma}_{1}^{\left(\mathrm{MNS}\right)}$ can be rewritten as
$\widehat{\gamma}_{1}^{\left(\mathrm{MNS}\right)}:={\displaystyle\sum\limits_{i=1}^{k}}a_{ik}\log\frac{Z_{n-i+1:n}}{Z_{n-k:n}},$
where
$a_{ik}:=\frac{\delta_{\left[n-i+1:n\right]}}{i}\prod_{j=i+1}^{k}\exp\left\{-\frac{\delta_{\left[n-j+1:n\right]}}{j}\right\}.$ (1.6)
This representation highlights the connection between the Nelson–Aalen weights $a_{ik}$ and the Kaplan–Meier-based integration, and makes explicit the recursive weighting scheme that adjusts for censoring in the upper tail.
The asymptotic properties of both the Kaplan-Meier-based estimator $\widehat{\gamma}_{1}^{\left(\mathrm{W}\right)}$ and the Nelson-Aalen-based estimator $\widehat{\gamma}_{1}^{\left(\mathrm{MNS}\right)}$ are established provided that $p>1/2.$ This condition ensures that a sufficiently large fraction of extreme observations remains uncensored, which is essential for reliable estimation of $\gamma_{1}.$ Equivalently, it can be expressed as $\gamma_{1}<\gamma_{2},$ meaning that the censoring distribution has a lighter tail than the distribution of interest. If $p\leq 1/2$ (or $\gamma_{1}\geq\gamma_{2}$), the effective tail sample size may be too small, and the Gaussian approximation underlying these estimators may break down. In practice, this condition serves as a natural guideline for the applicability of both estimators, emphasizing the importance of the relative tail heaviness between the observed and censoring distributions.
[…]
The most common method for estimating the parameters of an extreme value distribution in extreme value analysis relies on maximum likelihood estimation (MLE) method. Although these estimators possess desirable asymptotic properties, they can be sensitive to outlying observations from the assumed extreme value models; see, for instance, Brazauskas and Serfling (2000). Consequently, robust statistical methods provide a better alternative to mitigate the influence of outliers and deviations from the underlying parametric models. It has been shown that employing robust statistical ideas in extreme value theory improves the quality and precision of estimates (Dell’Aquila and Embrechts, 2006).
Juárèz and Schucany (2004) appear to be the first authors to employ the minimum density power divergence (MDPD) of Basu et al. (1998) for the robust estimation of the parameters of an extreme value distribution. Since then, this divergence measure has become a widely used tool for robust estimation of parameters of extreme value distributions. Kim and Lee (2008), Dierchx et al. (2013), Goegebeur et al. (2014), and Dierckx et al. (2021) have applied the MDPD approach to estimate the tail index and quantiles from Pareto-type distributions. Recently, Ghosh (2017) proposed a robust MDPD estimator for a real-valued tail index. This estimator is a robust generalization of the estimator proposed by Mattys and Beirlant (2003) and addresses non-identical distributions in the exponential regression model using the approach in Ghosh and Basu (2013). Furthermore, Dierchx et al. (2013) employed the MDPD concept on an extended Pareto distribution for relative excesses over a high threshold. Minkah el al. (2023A) and Minkah el al. (2023B) recently proposed a robust estimator for the tail index of a Pareto-type distribution using the MDPD approach applied to an exponential regression model.
While robust estimation techniques have been extensively developed for complete data settings, their extension to the context of randomly right-censored observations remains comparatively less explored. Such censoring mechanisms are prevalent in survival analysis, reliability theory, and actuarial applications, where incomplete information introduces additional statistical challenges. Classical estimators of the extreme value index under censoring, such as Hill-type approaches or those based on Kaplan-Meier and Nelson-Aalen integrals, often remain sensitive to outliers and model deviations, thereby limiting their practical robustness in real-world applications.
To the best of our knowledge, apart from the recent work of Dierckx et al. (2021), there exists no robust estimator specifically tailored to the tail index under random censoring. Their approach focuses on the conditional tail index, employing kernel smoothing techniques and local threshold selection for each covariate value. While the estimator effectively reduces bias and variance in the presence of contamination, it relies on a separate selection of tuning parameters and is restricted to the conditional setting, which complicates its direct application to the unconditional tail index. This limitation underscores the need for a robust and flexible framework capable of handling both unconditional and censored extreme value scenarios, motivating the development of the MDPD-based estimator proposed in this work.
The remainder of this paper is organized as follows. Section 2 provides a brief overview of the MDPD estimation method, originally introduced by Basu et al. (1998), highlighting its robust properties in the presence of outliers. Section 3 discusses in detail the asymptotic properties (consistency and asymptotic normality) of the proposed estimator, with full proofs deferred to Section 8. The finite-sample performance of the estimator is examined in Section 4, through a comprehensive simulation study, including comparisons with existing tail index estimators under censoring and contamination. Section 5 presents real-data applications, including a dataset on insurance claims (where censoring is relatively light, $p>1/2$) and the classical AIDS survival dataset, illustrating the practical relevance and robustness of the proposed methodology.
Finally, Appendix A compiles several technical lemmas and propositions that are instrumental for the theoretical development, while Appendix B contains the figures and additional results related to the simulation study.
2. Minimum density power divergence
Basu et al. (1998) introduced a new measure of divergence between two probability densities $\ell$ and $f,$ known as the density power divergence (DPD), defined by
$d_{\alpha}\left(f,\ell\right):=\left\{\begin{array}[c]{ll}\int_{\mathbb{R}}\left[\ell^{1+\alpha}\left(x\right)-\left(1+\dfrac{1}{\alpha}\right)\ell^{\alpha}\left(x\right)f\left(x\right)+\dfrac{1}{\alpha}f^{1+\alpha}\left(x\right)\right]dx,&\alpha>0\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \int_{\mathbb{R}}f\left(x\right)\log\dfrac{f\left(x\right)}{\ell\left(x\right)}dx,&\alpha=0,\end{array}\right$. (2.7)
where $\alpha\geq 0$ is a tuning parameter controlling the trade-off between efficiency and robustness. The case $\alpha=0$ is the limit of the general expression $\left(\alpha>0\right)$ as $\alpha\downarrow 0$ yielding the classical Kullback-Leibler divergence $d_{0}\left(f,\ell\right).$ Let us consider a parametric family of densities $\left\{\ell_{\theta}:\Theta\subset\mathbb{R}^{p}\right\}$ and suppose that the goal is to estimate the parameter $\theta.$ Let $F$ denote the cdf corresponding to the density $f.$ The MDPD functional $T_{\alpha}\left(F\right)$ is defined as
$d_{\alpha}\left(f,\ell_{T_{\alpha}\left(F\right)}\right)=\min_{\theta\in\Theta}d_{\alpha}\left(f,\ell_{\theta}\right).$
[…]
$\delta_{\alpha}\left(f;\theta\right):=\left\{\begin{array}[c]{lc}\int_{\mathbb{R}}\ell_{\theta}^{1+\alpha}\left(x\right)dx-\left(1+\dfrac{1}{\alpha}\right)\int_{\mathbb{R}}\ell_{\theta}^{\alpha}\left(x\right)dF\left(x\right),&\alpha>0,\vskip 6.0pt plus 2.0pt minus 2.0pt\\ -\int_{\mathbb{R}}\log\ell_{\theta}\left(x\right)dF\left(x\right),&\alpha=0.\end{array}\right$.
Given a sample $X_{1},...,X_{n}$ from the cdf $F,$ we estimate $\delta_{\alpha}\left(f;\theta\right)$ by replacing $F$ with its empirical counterpart $F_{n}\left(x\right):=n^{-1}\sum_{i=1}^{n}\mathbb{I}_{\left\{X_{i}\leq x\right\}}.$ The MDPD estimator is then the minimizer (over $\theta\in\Theta)$ of the empirical objective function
$L_{n.\alpha}\left(\theta\right)=\left\{\begin{array}[c]{lc}\int_{\mathbb{R}}\ell_{\theta}^{1+\alpha}\left(x\right)dx-\left(1+\dfrac{1}{\alpha}\right)\dfrac{1}{n}{\displaystyle\sum\limits_{i=1}^{n}}\ell_{\theta}^{\alpha}\left(X_{i}\right),&\alpha>0,\vskip 6.0pt plus 2.0pt minus 2.0pt\\ -\dfrac{1}{n}{\displaystyle\sum\limits_{i=1}^{n}}\log\ell_{\theta}\left(X_{i}\right),&\alpha=0.\end{array}\right.$.
That is, $\widehat{\theta}_{n,\alpha}:=\arg\min_{\theta\in\Theta}L_{n.\alpha}\left(\theta\right),$ which satisfies the estimating equation
$\left\{\begin{array}[c]{lc}\int_{\mathbb{R}}\dfrac{d}{d\theta}\ell_{\theta}^{1+\alpha}\left(x\right)dx-\left(1+\dfrac{1}{\alpha}\right)\dfrac{1}{n}{\displaystyle\sum\limits_{i=1}^{n}}\dfrac{d}{d\theta}\ell_{\theta}^{\alpha}\left(X_{i}\right)=0,&\alpha>0,\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \dfrac{1}{n}{\displaystyle\sum\limits_{i=1}^{n}}\dfrac{d}{d\theta}\log\ell_{\theta}\left(X_{i}\right)=0,&\alpha=0.\end{array}\right$.
Small $\alpha$ yields high efficiency with limited robustness, while moderate $\alpha$ values provide strong robustness under contamination with minor efficiency loss compared to MLE under correct model specification.
2.1. MDPD estimator of $\gamma_{1}$ for right-censored data
For a fixed threshold $u>0,$ consider the relative excess rv $Y:=X/u$ conditional on $X>u.$ Its cdf is given by
$F_{u}\left(x\right):=1-\frac{\overline{F}\left(ux\right)}{\overline{F}\left(u\right)},$
with corresponding probability density function $f_{u}.$ Assuming that $\overline{F}\in\mathcal{RV}_{\left(-1/\gamma_{1}\right)},$ it follows that $F_{u}\left(x\right)\approx 1-x^{-1/\gamma_{1}},$ as $u\rightarrow\infty.$ Therefore, the rv $Y$ may be approximated by the Pareto distribution with density
$\ell_{\gamma_{1}}\left(x\right):=\frac{d}{dx}\left(1-x^{-1/\gamma_{1}}\right)=\gamma_{1}^{-1}x^{-1-1/\gamma_{1}},\text{ for }x\geq 1.$
The corresponding DPD objective function is
$\mathcal{L}_{u,\alpha}^{\ast}\left(\gamma_{1}\right):=\left\{\begin{array}[c]{lc}\int_{1}^{\infty}\ell_{\gamma_{1}}^{1+\alpha}\left(x\right)dx-\left(1+\dfrac{1}{\alpha}\right)\int_{u}^{\infty}\ell_{\gamma_{1}}^{\alpha}\left(x/u\right)d\dfrac{F\left(x\right)}{\overline{F}\left(u\right)},&\alpha>0,\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \int_{u}^{\infty}\log\ell_{\gamma_{1}}\left(x/u\right)d\dfrac{F\left(x\right)}{\overline{F}\left(u\right)},&\alpha=0,\end{array}\right$. (2.8)
Letting $u=Z_{n-k:n}$ and replacing $F$ with the Nelson-Aalen estimator $F_{n}^{\left(NA\right)}$ gives the empirical objective function
$\mathcal{L}_{k,\alpha}^{\ast}\left(\gamma_{1}\right):=\left\{\begin{array}[c]{lc}\int_{1}^{\infty}\ell_{\gamma_{1}}^{1+\alpha}\left(x\right)dx-\left(1+\dfrac{1}{\alpha}\right)\int_{1}^{\infty}\ell_{\gamma_{1}}^{\alpha}\left(x/Z_{n-k:n}\right)d\dfrac{F_{n}^{\left(NA\right)}\left(x\right)}{\overline{F}_{n}^{\left(NA\right)}\left(Z_{n-k:n}\right)},&\alpha>0,\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \int_{1}^{\infty}\log\ell_{\gamma_{1}}\left(x/Z_{n-k:n}\right)d\dfrac{F_{n}^{\left(NA\right)}\left(x\right)}{\overline{F}_{n}^{\left(NA\right)}\left(Z_{n-k:n}\right)},&\alpha=0.\end{array}\right$. (2.9)
The MDPD tail index estimator under random censoring, denoted $\widehat{\gamma}_{1,\alpha},$ is then obtained by minimizing $\mathcal{L}_{k,\alpha}^{\ast}\left(\gamma_{1}\right).$ That is, by solving the following equation:
$\left\{\begin{array}[c]{lc}\begin{array}[c]{l}\int_{1}^{\infty}\dfrac{d}{d\gamma_{1}}\ell_{\gamma_{1}}^{1+\alpha}\left(\dfrac{x}{Z_{n-k:n}}\right)dx\\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -\left(1+\dfrac{1}{\alpha}\right)\int_{1}^{\infty}\dfrac{d}{d\gamma_{1}}\ell_{\gamma_{1}}^{\alpha}\left(\dfrac{x}{Z_{n-k:n}}\right)d\dfrac{F_{n}^{\left(NA\right)}\left(x\right)}{\overline{F}_{n}^{\left(NA\right)}\left(Z_{n-k:n}\right)}=0,\end{array}&\alpha>0,\vskip 6.0pt plus 2.0pt minus 2.0pt\\ \int_{1}^{\infty}\dfrac{d}{d\gamma_{1}}\log\ell_{\gamma_{1}}\left(\dfrac{x}{Z_{n-k:n}}\right)d\dfrac{F_{n}^{\left(NA\right)}\left(x\right)}{\overline{F}_{n}^{\left(NA\right)}\left(Z_{n-k:n}\right)}=0,&\alpha=0.\end{array}\right$.
By elementary calculations:
$\dfrac{d}{d\gamma_{1}}\ell_{\gamma_{1}}^{\alpha}\left(x\right)=\frac{\alpha}{\gamma_{1}^{2+\alpha}}\frac{\log x-\gamma_{1}}{x^{\alpha\left(1+1/\gamma_{1}\right)}},$
$\dfrac{d}{d\gamma_{1}}\log\ell_{\gamma_{1}}\left(x\right)=\dfrac{d}{d\gamma_{1}}\log\left(\gamma_{1}^{-1}x^{-1-1/\gamma_{1}}\right)=-\frac{\gamma_{1}-\log x}{\gamma_{1}^{2}}$
and
$\int_{1}^{\infty}\dfrac{d}{d\gamma_{1}}\ell_{\gamma_{1}}^{1+\alpha}\left(x\right)dx=-\frac{\alpha\left(\alpha+1\right)\left(\gamma_{1}+1\right)}{\gamma_{1}^{\alpha+1}\left(\alpha\left(1+\gamma_{1}\right)+1\right)^{2}}.$
Hence, the estimating equation for $\widehat{\gamma}_{1,\alpha}$ is
$\left\{\begin{array}[c]{lc}{\displaystyle\int_{Z_{n-k:n}}^{\infty}}\dfrac{\gamma_{1}-\log\left(x/Z_{n-k:n}\right)}{\left(x/Z_{n-k:n}\right)^{\alpha\left(1+1/\gamma_{1}\right)}}d\dfrac{F_{n}^{\left(NA\right)}\left(x\right)}{\overline{F}_{n}^{\left(NA\right)}\left(Z_{n-k:n}\right)}=\dfrac{\alpha\gamma_{1}\left(\gamma_{1}+1\right)}{\left(1+\alpha+\alpha\gamma_{1}\right)^{2}},&\alpha>0,\vskip 6.0pt plus 2.0pt minus 2.0pt\\ {\displaystyle\int_{Z_{n-k:n}}^{\infty}}\left(\gamma_{1}-\log\dfrac{x}{Z_{n-k:n}}\right)d\dfrac{F_{n}^{\left(NA\right)}\left(x\right)}{\overline{F}_{n}^{\left(NA\right)}\left(Z_{n-k:n}\right)}=0,&\alpha=0.\end{array}\right$. (2.10)
Meraghni et al. (2025) stated that
$dF_{n}^{\left(\mathrm{NA}\right)}\left(x\right)=\frac{\overline{F}_{n}^{\left(\mathrm{NA}\right)}\left(x\right)dH_{n}^{\left(1\right)}\left(x\right)}{\overline{H}_{n}\left(x^{-}\right)},$ (2.11)
where $H_{n}$ and $H_{n}^{\left(1\right)}$ are the empirical counterparts of the cdf $H$ and the sub-distribution $H^{\left(1\right)},$ respectively defined by:
$H_{n}\left(x\right):=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}_{\left\{Z_{i:n}\leq x\right\}}\text{ and }H_{n}^{\left(1\right)}\left(x\right):=\frac{1}{n}\sum_{i=1}^{n}\delta_{\left[i:n\right]}\mathbb{I}_{\left\{Z_{i:n}\leq x\right\}},$
[…]
$\left\{\begin{array}[c]{lc}{\displaystyle\sum\limits_{i=1}^{k}}a_{ik}\dfrac{\gamma_{1}-\log\left(Z_{n-i+1:n}/Z_{n-k:n}\right)}{\left(Z_{n-i+1:n}/Z_{n-k:n}\right)^{\alpha\left(1+1/\gamma_{1}\right)}}=\dfrac{\alpha\gamma_{1}\left(\gamma_{1}+1\right)}{\left(1+\alpha+\alpha\gamma_{1}\right)^{2}},&\alpha>0,\vskip 6.0pt plus 2.0pt minus 2.0pt\\ {\displaystyle\sum\limits_{i=1}^{k}}a_{ik}\left\{\gamma_{1}-\log\dfrac{Z_{n-i+1:n}}{Z_{n-k:n}}\right\}=0,&\alpha=0,\end{array}\right$. (2.12)
[…]
This estimator provides the first robust generalization of the tail index under random censoring using the MDPD approach, bridging the gap between efficiency and robustness in extreme value estimation.
3. Consistency and asymptotic normality
From now on, we denote the true value of $\gamma_{1}$ by $\gamma_{1}^{\ast}.$
Theorem 3.1.
[…]
$k\rightarrow\infty\text{ and }k/n\rightarrow 0,\text{ as }n\rightarrow\infty.$
[…]
$\widehat{\gamma}_{1,\alpha}\overset{\mathbf{P}}{\rightarrow}\gamma_{1}^{\ast},\text{ as }n\rightarrow\infty.$
This result establishes the consistency of the MDPD tail index estimator under random censoring, underlining that a sufficient proportion of uncensored extreme observations $(p>1/2)$ is crucial for reliable estimation.
Since weak approximations of extreme value theory-based statistics are typically achieved within the second-order framework (see, i.e., de Haan and Stadtmüller, 1996), we assume that the cdf $F$ satisfies the well-known second-order condition of regular variation. Specifically, for any $x>0:$
$\underset{t\rightarrow\infty}{\lim}\frac{U_{F}\left(tx\right)/U_{F}\left(t\right)-x^{\gamma_{1}}}{A_{1}^{\ast}\left(t\right)}=x^{\gamma_{1}}\dfrac{x^{\tau_{1}}-1}{\tau_{1}},$ (3.13)
where $\tau_{1}\leq 0$ is the second-order parameter, and $A_{1}^{\ast}$ is a function tending to $0,$ not changing sign near infinity and whose absolute value is regularly varying with index $\tau_{1}.$ When $\tau_{1}=0,$ the expression $\left(x^{\tau_{1}}-1\right)/\tau_{1}$ is interpreted as $\log x.\vskip 3.0pt plus 1.0pt minus 1.0pt$
We denote the quantile and tail quantile functions of a given cdf $D$ by
$\mathcal{D}^{\leftarrow}\left(s\right):=\inf\left\{x:\mathcal{D}\left(x\right)\geq s\right\},\text{ }0<s<1$
and
$U_{\mathcal{D}}\left(t\right):=\mathcal{D}^{\leftarrow}\left(1-1/t\right),t>1,$
respectively.
For convenience, we set $h:=U_{H}\left(n/k\right)$ and define $A_{1}\left(t\right):=A_{1}^{\ast}\left(1/\overline{F}\left(t\right)\right),$ for $t>1.$
Theorem 3.2.
[…]
Then for any $\alpha>0:$
$\left(1+\frac{1}{\alpha}\right)^{-1}\eta^{\ast}\sqrt{k}\left(\widehat{\gamma}_{1,\alpha}-\gamma_{1}^{\ast}\right)\overset{\mathcal{D}}{\rightarrow}\mathcal{N}\left(\lambda\mu,\sigma^{2}\right),\text{ as }n\rightarrow\infty,$
[…]
This theorem establishes the asymptotic normality of the MDPD estimator under random censoring, emphasizing again that the condition $p>1/2$ is essential for the validity of the Gaussian approximation.
4. Simulation Study
In this section, we investigate the finite-sample performance of the proposed MDPD tail index estimator $\widehat{\gamma}_{1,\alpha}$ through an extensive Monte Carlo simulation study. The main objective is to assess both efficiency and robustness in the presence of random right censoring and contamination affecting the upper tail of the distribution. Particular attention is paid to the impact of contamination on bias, MSE, and stability with respect to the choice of the number of upper order statistics. The simulation design is chosen so as to remain fully consistent with the theoretical framework developed in the previous sections.
4.1. Compared estimators and efficiency–robustness trade-off
The MDPD estimator $\widehat{\gamma}_{1,\alpha}$ is computed for $\alpha\in\{0.1,0.3,0.5\},$ allowing us to explore the efficiency–robustness trade-off induced by the tuning parameter. Small values of $\alpha$ emphasize efficiency under correct model specification, whereas larger values enhance robustness against contamination in the upper tail.
For comparison purposes, the following benchmark estimators are included:
  • the adapted Hill-type estimator $\widehat{\gamma}_{1}^{(\mathrm{EFG})}$,
  • the Kaplan–Meier-based estimator $\widehat{\gamma}_{1}^{(\mathrm{W})},$
  • the Nelson–Aalen-based estimator $\widehat{\gamma}_{1}^{(\mathrm{MNS})},$ corresponding to the special case $\widehat{\gamma}_{1,0}.$
This selection encompasses both classical and censoring-adapted tail index estimators that are standard in the extreme value literature. It therefore provides a relevant and informative benchmark for assessing the efficiency and robustness of the proposed MDPD-based estimator under random censoring.
4.2. Simulation models and censoring design
We consider two classical heavy-tailed distributions for the latent lifetime variable $X,$ which are widely used as benchmark models in extreme value analysis.
The first model is the Burr distribution, defined by
$F(x)=1-\left(1+x^{1/\eta}\right)^{-\eta/\gamma_{1}},\qquad x>0,$
with shape parameter $\eta=0.25$ and tail index $\gamma_{1}\in\{0.3,0.5\}.$
The second model is the Fréchet distribution, defined by
$F(x)=\exp\left(-x^{-1/\gamma_{1}}\right),\qquad x>0,$
with tail index $\gamma_{1}\in\{0.3,0.5\}.$
Right censoring is generated independently through a random variable $C$ drawn from the same parametric family as $X,$ with censoring tail index $\gamma_{2}.$ The model parameters are chosen such that the proportion of uncensored extreme observations satisfies $p>1/2.$ This condition ensures that the observed variable $Z=\min(X,C)$ remains regularly varying and fulfills the weak censoring condition introduced in Section 1. This construction allows us to control the censoring mechanism while preserving the asymptotic tail behavior of the observed data.
The choices of the censoring level and the contamination proportion are guided by both theoretical requirements and practical considerations. The proportion of uncensored extreme observations is fixed at $p\in\{0.55,0.7\},$ corresponding respectively to moderate and mild censoring regimes. Both values satisfy the weak censoring condition $p>1/2,$ which guarantees identifiability of the tail index and preserves regular variation of the observed data. Varying $p$ allows us to assess the impact of censoring severity on estimation accuracy while remaining within a statistically coherent framework.
4.3. Contamination schemes
To assess robustness, we consider contamination scenarios of increasing severity, ranging from the ideal uncontaminated setting to adverse configurations in which contamination affects the upper tail of the lifetime distribution. The contamination intensity is controlled by a parameter $\epsilon\in(0,1),$ which is common to all contaminated configurations.
In all contaminated settings, contamination is modeled through a mixture mechanism acting on the latent lifetime variable. Specifically, the distribution of $X$ is given by
$F_{\epsilon}=(1-\epsilon)F+\epsilon F_{c},$
where $F$ denotes the target distribution and $F_{c}$ is a heavier-tailed contaminating distribution with tail index $\gamma_{c}>\gamma_{1}.$ This mixture formulation offers a flexible and interpretable framework for robustness analysis, capturing the presence of extreme outliers while maintaining identifiability under independent censoring.
The condition $\gamma_{c}>\gamma_{1}$ ensures that contamination primarily affects the upper tail of the distribution, generating extreme observations that act as outliers relative to the target model. This setting reflects realistic departures from the ideal model and provides a stringent stress test for robustness. The resulting mixture formulation offers a flexible and interpretable framework for robustness analysis while preserving identifiability under independent censoring.
Several contamination configurations are considered:
In the uncontaminated case $(\epsilon=0),$ the lifetime variable $X$ and the censoring variable $C$ are generated independently from their target distributions $F$ and $G,$ respectively. The observed sample $(Z,\delta)$ is obtained through the standard random right-censoring mechanism. This configuration serves as a benchmark for assessing efficiency under correct model specification and for quantifying the efficiency loss induced by robustness-oriented procedures.
In the pre-censoring contamination setting, contamination is introduced at the level of the latent lifetime variable $X,$ prior to censoring. A proportion $\epsilon\in(0,1)$ of the $X$-values is replaced by extreme observations generated from a heavier-tailed distribution with tail index $\gamma_{c}>\gamma_{1}.$ The censoring variable $C$ is generated independently from its target distribution $G,$ whose tail behavior is governed by $\gamma_{2}.$ The censoring mechanism is then applied to the contaminated latent sample, yielding the observed pairs $(Z,\delta).$ This framework preserves the independent censoring structure and directly reflects the impact of contamination on the tail behavior of the lifetime distribution, making it a natural, coherent, and identifiable setting for robustness analysis in both parametric and nonparametric frameworks.
An alternative contamination mechanism consists in perturbing the observable pair $(Z,\delta)$ after the censoring mechanism has taken place. While such post-censoring contamination may be interpreted as data corruption or recording errors at the observational stage, it does not correspond to a well-defined censoring model in a nonparametric framework. By directly altering the joint distribution of $(Z,\delta),$ this mechanism breaks the link between the observed data and the underlying lifetime and censoring distributions $(F,G),$ thereby rendering reconstruction-based inference ill-posed. For this reason, post-censoring contamination is not retained in the simulation study and is viewed only as an illustrative stress mechanism rather than as a statistically coherent robustness framework.
Finally, contamination affecting the censoring variable $C$ is not considered. Such contamination would alter the censoring mechanism itself and may violate the independent censoring assumption underlying most theoretical results for censored extreme value estimation. In particular, contamination of $C$ directly modifies the joint distribution of the observable pair $(Z,\delta)$ in a way that cannot generally be associated with any well-defined latent lifetime distribution $F$. As a consequence, tail index estimation loses a clear statistical interpretation. Accordingly, robustness assessments in this paper focus exclusively on contamination acting on the lifetime variable prior to censoring.
4.4. Simulation scenarios and settings
Although infinitely many combinations of target, censoring, and contamination distributions could be considered, we restrict attention to two representative scenarios. These scenarios are standard in the extreme value literature and allow for a clear and informative assessment of both efficiency and robustness under censoring and contamination.
In Scenario S1 (Burr/Burr/Burr), the target variable $X$ follows a Burr distribution, the censoring variable $C$ is also Burr-distributed, and contamination is introduced through a Burr distribution with parameters $\eta_{c}=0.25$ and $\gamma_{c}\in\{0.6,0.8\}.$
In Scenario S2 (Fréchet/Fréchet/Fréchet), the target variable $X$ follows a Fréchet distribution, the censoring mechanism is Fréchet-distributed, and contamination is generated from a Fréchet distribution with tail index $\gamma_{c}\in\{0.6,0.8\}.$
These configurations allow us to examine robustness across different tail behaviors while maintaining a coherent and interpretable simulation framework. Contamination intensity is controlled by the parameter $\epsilon\in(0,1),$ which governs the proportion of observations affected by contamination in the upper tail.
For each scenario, contamination levels $\epsilon\in\{0,0.15,0.40\}$ are considered. These values correspond respectively to an uncontaminated setting, a moderately contaminated setting, and a severely contaminated configuration designed to stress-test robustness properties while remaining within a controlled and interpretable simulation framework.
4.5. Simulation results under pre-censoring contamination
This section investigates the finite-sample performance of the proposed MDPD-based tail index estimator under pre-censoring contamination, that is, when contamination affects the latent lifetime variable prior to the censoring mechanism. As emphasized in Section 4, this contamination framework is statistically coherent with both parametric and nonparametric censored models, preserves the independent censoring structure, and provides a meaningful benchmark for assessing robustness with respect to perturbations of the underlying tail distribution.
Scenario S1 (Burr distribution).
Case $\gamma_{1}=0.3$ (Figures 10.110.2). In the absence of contamination $\left(\epsilon=0\right),$ all estimators exhibit comparable behavior for intermediate values of the number of upper order statistics $k.$ As expected, classical estimators such as $\widehat{\gamma}_{1}^{(\mathrm{EFG})},$ $\widehat{\gamma}_{1}^{(\mathrm{W})},$ and $\widehat{\gamma}_{1}^{(\mathrm{MNS})}$ achieve good efficiency in this ideal setting, although their performance remains sensitive to the choice of the threshold parameter $k,$ a well-known feature of tail index estimation.
When contamination is introduced, a markedly different behavior emerges. Classical estimators deteriorate rapidly as $k$ increases, displaying pronounced bias and instability even for moderate contamination levels. This degradation reflects their intrinsic sensitivity to extreme observations generated by the heavier-tailed contaminating component. In contrast, the MDPD estimators exhibit a clear robustness effect: while small values of $\alpha$ retain some sensitivity to contamination, the estimators corresponding to $\alpha\geq 0.3$ remain stable over a wide range of $k,$ with substantially reduced bias and MSE. This clearly illustrates the role of the tuning parameter $\alpha$ in controlling the influence of outlying extremes and stabilizing inference.
Case $\gamma_{1}=0.5$ (Figures 10.310.4). For heavier tails, the adverse effect of contamination becomes more pronounced. Even in the uncontaminated case, classical estimators show increased variability, reflecting the intrinsic difficulty of tail index estimation in heavier-tailed regimes. Under contamination, these estimators become severely biased and unstable, with performance degrading rapidly as $k$ grows.
By contrast, the MDPD estimators maintain a remarkable degree of stability. In particular, estimators with $\alpha\geq 0.3$ exhibit nearly flat trajectories across a broad range of $k,$ indicating strong robustness with respect to both tail heaviness and contamination intensity. These results confirm that robustification becomes increasingly critical as the tail index increases, that is, when extremes play a dominant role.
Scenario S2 (Fréchet distribution).
Case $\gamma_{1}=0.3$ (Figures 10.510.6). The qualitative behavior observed in the Burr case carries over to the Fréchet setting. In the absence of contamination, all estimators perform satisfactorily for moderate values of $k,$ although classical estimators again display noticeable sensitivity to the choice of the threshold.
Under contamination, classical estimators experience a substantial loss of accuracy, with increasing bias and variability. This confirms that their lack of robustness is not distribution-specific but persists across different heavy-tailed models. In contrast, the MDPD estimators demonstrate a robust behavior similar to that observed in Scenario S1, with larger values of $\alpha$ yielding improved stability and resistance to contamination-induced distortions.
Case $\gamma_{1}=0.5$ (Figures 10.710.8). This configuration represents the most adverse scenario, combining heavy tails with severe contamination. In this setting, classical estimators perform particularly poorly, often exhibiting dramatic bias and erratic behavior across the range of $k$ values.
The MDPD estimators, however, remain reliable even in this challenging regime. Although a moderate increase in variance is observed for very large values of $\alpha,$ this effect is largely compensated by a substantial reduction in bias, so that the overall MSE remains significantly lower than that of the classical competitors. This highlights the practical advantage of the MDPD approach in extreme scenarios where standard methods effectively break down.
Overall discussion.
Overall, the simulation results provide strong empirical evidence that the proposed MDPD-based tail index estimator offers a robust and reliable alternative for censored extreme value analysis. While classical estimators may retain efficiency under ideal conditions, their performance deteriorates sharply in the presence of contaminated extremes. In contrast, the MDPD estimators achieve a favorable efficiency–robustness trade-off, particularly for moderate to large values of the tuning parameter $\alpha.$
Taken together, these findings demonstrate that the proposed methodology is well suited for realistic applications, where both censoring and contamination are unavoidable and robustness with respect to extreme observations is a key requirement.
5. Real data application
5.1. Insurance loss data (weak censoring)
We applied our estimation procedures to an insurance loss dataset available in the R package copula, originally collected by the Insurance Services Office, Inc. The dataset contains $1500$ observations, among which $34$ are right-censored because the actual loss exceeds the policy limit, which varies across contracts. The observed variables are:
  • $X_{j}:$ actual loss amount of claim $j.$
  • $C_{j}:$ policy limit for contract $j=1,\dots,1500.$
Because of policy limits, some losses are censored when $X_{j}>C_{j}.$ In these cases, the observed value is
$Z_{j}=\min(X_{j},C_{j}),$
with censoring indicator
$\delta_{j}=\mathbb{I}_{\{X_{j}\leq C_{j}\}}.$
These censored claims correspond to contracts where reported losses reach the policy ceiling, indicating that the true loss exceeds the observed value. They reflect the presence of extreme losses in the upper tail but do not fully capture all extreme observations. This dataset has been examined in prior studies, such as Frees and Valdez (1998), Klugman and Parsa (1999), and Denuit et al. (2006).
Before proceeding to tail index estimation, it is essential to assess the effective amount of tail information available in the presence of censoring. In particular, we first estimate the proportion of uncensored extreme observations in order to verify that the weak censoring assumption required by the proposed estimator is satisfied.
5.2. Estimation of the proportion $p$ of non-censored extremes
To justify the application of the MDPD estimator under weak censoring, we estimate the proportion $p$ of uncensored extreme observations. This quantity plays a central role, as it determines whether a sufficient number of extreme observations remain fully observed to support reliable tail inference.
The adaptive algorithm of Reiss and Thomas is applied to select the optimal number of upper order statistics $k^{\ast}$ used for estimating $p$. This data-driven procedure allows us to balance stability and sensitivity in the tail region, while accounting for the presence of censoring.
5.3. Selection of the number of upper order statistics
The number of upper order statistics $k$ is selected using the adaptive algorithm of Reiss and Thomas Reiss and Thomas (2007), which aims at stabilizing tail index estimation based on uncensored extremes. Specifically, the optimal value $k^{\ast}$ is defined as
$k^{\ast}:=\arg\min_{1<k<n}\frac{1}{k}\sum_{i=1}^{k}i^{\theta}\left|\widehat{\zeta}_{i}-\text{median}(\widehat{\zeta}_{1},\dots,\widehat{\zeta}_{k})\right|,\text{ }0\leq\theta\leq 0.5,$
where $\widehat{\zeta}_{i}$ denotes an estimator of the tail parameter $\eta$ computed from the $i$ largest order statistics.
Throughout this work, we fix $\theta=0.3.$ This choice is supported by an extensive empirical calibration based on a preliminary study of $\theta$ over the interval $[0,$ $0.5],$ which systematically examined the sensitivity of the procedure to the tuning parameter and showed that $\theta=0.3$ consistently provides a favorable compromise in terms of bias and mean squared error across the considered scenarios. Smaller values of $\theta$ tend to oversmooth the variability among the upper order statistics, thereby masking relevant tail information, whereas larger values overemphasize extreme fluctuations, resulting in increased variance and reduced numerical stability.
Thus, fixing $\theta=0.3$ offers a stable and numerically well-justified balance between robustness and sensitivity to extremes. This value was found to perform reliably across different tail indices, censoring levels, and contamination regimes. The Reiss–Thomas selection procedure is applied uniformly throughout the simulation study, thereby ensuring a fair comparison between the different estimators. Further details on this selection rule can be found in Neves and Fraga Alves (2004).
Applying this procedure to the insurance loss data yields an optimal value $k^{\ast}=51$ for estimating $p.$ The resulting estimate $\widehat{p}=0.76,$ confirms that the weak censoring condition $(p>1/2)$ holds in this dataset, validating the applicability of the proposed MDPD-based approach.
Note. In this real-data analysis, we do not introduce artificial contamination. The MDPD estimator is already robust both from a theoretical perspective and as confirmed by the simulation results. Moreover, there is no clear empirical evidence suggesting contamination in the insurance data, and artificially modifying the observations could distort the underlying loss distribution. The dataset is therefore analyzed as observed, reflecting realistic insurance losses, including naturally occurring extreme values.
5.4. Tail index estimation using MDPD
Having verified the weak censoring assumption and determined an appropriate threshold, we now proceed to tail index estimation. The same Reiss–Thomas algorithm is employed to select the number of upper order statistics for tail index estimation, leading to $k^{\ast}=73,$ This value is kept fixed across all robustness settings to facilitate comparison.
The MDPD estimator is computed for four robustness parameter values:
$\alpha\in\{0.01,0.1,0.3,0.5\},$
allowing us to assess the impact of robustness on tail index estimation in a realistic insurance context.
5.5. Contamination scenarios
For real data, we report the analysis on the observed dataset without artificial contamination.
  • The original, observed data are treated as the reference scenario.
    Note. In contrast to simulations, contamination is not applied to real data, because:
  • The MDPD estimator is theoretically robust and performs well under heavy-tailed observations.
  • There is no confirmed contamination in the insurance dataset, and applying artificial extreme values could distort the estimation.
  • Weak censoring $(p>1/2)$ is satisfied, allowing meaningful application of the estimator.
By focusing on the actual observed losses, the effect of the robustness parameter $\alpha$ can still be assessed, showing how the estimator adapts to potentially extreme claims in the dataset.
5.6. Estimated tail indices
The estimated tail indices for different values of $\alpha$ are summarized in Table 5.1. Increasing $\alpha$ provides robustness against extreme values that may naturally occur in the dataset without artificially contaminating the data.
$\alpha$ $\widehat{\gamma}_{1,\alpha}$
$0.01$ $0.745$
$0.1$ $0.773$
$0.3$ $0.820$
$0.5$ $0.845$
Table 5.1. MDPD estimates of the tail index $\gamma_{1}$ for insurance loss data
5.7. Discussion and key takeaways
  • Small $\alpha$ values (0.010.1)(0.01-0.1) produce estimates close to classical estimators, maintaining high efficiency for the bulk of the data, while larger $\alpha$ (0.30.5)(0.3-0.5) slightly overestimate the tail index but improve robustness against potential outliers.
  • Larger $\alpha$ values (0.30.5)(0.3-0.5) slightly increase the estimated tail index, providing additional protection against extreme claims, while preserving interpretability.
[…]
In conclusion, Section 5 highlights that the MDPD approach can be applied directly to real datasets, capturing extreme events effectively while retaining robustness, efficiency, and interpretability.
5.8. Australian AIDS survival data (strong censoring)
The Australian AIDS survival dataset contains information on patients diagnosed with AIDS in Australia prior to July 1, 1991. The data were provided by Dr. P. J. Solomon and the Australian National Centre in HIV Epidemiology and Clinical Research. We focus on $2754$ male patients out of a total of $2843,$ consistent with prior studies such as Ripley and Solomon (1994) and Venables and Ripley (2002) (pages 379-385), Einmahl et al. (2008), Ndao et al. (2014) and Stupfler (2016).
The proportion of uncensored observations, denoted $p,$ is estimated using the Reiss–Thomas adaptive algorithm. The optimal number of upper order statistics for estimating $p$ is $k^{\ast}=162,$ yielding an estimated $\widehat{p}=0.29.$ This indicates strong right censoring, with less than one-third of the observations being uncensored.
Under such strong censoring, the theoretical assumptions for the MDPD tail index estimator $(p>1/2)$ are not satisfied. Consequently, the estimator’s performance may degrade, showing increased variability and potential bias. Attempts to compute the tail index using the Kaplan–Meier-based Worms–Worms estimator ($\widehat{\gamma}_{1}^{(\mathrm{W})}$) or the Nelson–Aalen-based MDPD estimator with $\alpha=0$ ($\gamma_{1}^{(\mathrm{MNS})}$) either fail to converge or produce unreliable results. Only the adapted Hill-type estimator ($\widehat{\gamma}_{1}^{(\mathrm{EFG})}$) provides estimates; however, it lacks robustness, inheriting the classical Hill estimator’s sensitivity to extreme observations.
Nevertheless, this dataset provides a valuable case for exploring the limits of the MDPD approach. It illustrates the challenges of estimating tail indices under strong censoring and highlights the need for methods capable of handling situations where the proportion of uncensored extremes is below $50\%$.
Overall, the real data analyses complement the simulation study and provide additional insight into the behavior of tail index estimators under censoring:
  • In the insurance dataset, which satisfies the weak censoring condition $(p>1/2),$ the proposed MDPD estimator performs according to theoretical expectations, exhibiting clear robustness gains when contamination affects the latent losses before censoring. The additional experiment applying contamination after censoring shows that such a mechanism does not correspond to a meaningful perturbation of the target distribution and may distort inference, justifying its illustrative-only use.
  • In contrast, the Australian AIDS survival dataset represents a strong censoring scenario $(p<1/2),$ where the assumptions underlying the MDPD estimator are violated. In this setting, most estimators designed for censored extremes fail to provide stable or reliable estimates. Even the adapted Hill-type estimator, although computable, remains non-robust due to its sensitivity to extreme observations.
These findings delineate the practical scope of the proposed methodology. The MDPD estimator provides an effective and robust solution for tail index estimation under weak censoring, which encompasses many applications in insurance and reliability analysis. At the same time, the strong censoring case, exemplified by the AIDS dataset, highlights open challenges and motivates future research on methodological extensions capable of handling extreme value inference beyond the $p>1/2$ regime.
6. Conclusion
In this work, we addressed the problem of robust estimation of the extreme value index under random right censoring, a framework that naturally arises in survival analysis, reliability engineering, and insurance applications. While a variety of estimators have been proposed in the literature—including Hill-type procedures and methods based on Kaplan–Meier or Nelson–Aalen estimators—many of them remain sensitive to outliers or rely on restrictive modeling assumptions, particularly in the presence of contamination.
To overcome these limitations, we introduced a new tail index estimator based on the Minimum Density Power Divergence (MDPD), a criterion well known for its robustness properties in uncensored settings. We carefully extended the MDPD methodology to the random right-censoring framework, explicitly accounting for the incomplete nature of the observed data. To the best of our knowledge, this work constitutes the first application of the MDPD framework to tail index estimation under random censoring, thereby filling an important gap in the robust extreme value literature.
Under standard regular variation assumptions and within the weak censoring regime, we established the consistency and asymptotic normality of the proposed estimator. A comprehensive simulation study confirmed its strong finite-sample performance, highlighting a favorable robustness–efficiency trade-off, especially under model misspecification and contamination.
The practical relevance of the approach was illustrated using an insurance loss dataset satisfying the weak censoring condition $(p>1/2).$ In this setting, the estimator behaves in accordance with the theoretical results and exhibits clear robustness gains when contamination affects the latent losses prior to censoring. In contrast, the analysis of the Australian AIDS survival dataset, characterized by strong censoring $(p\leq 1/2),$ underscores the intrinsic difficulties of tail inference in such regimes, where most existing estimators fail to deliver stable or reliable results. This empirical evidence also confirms that contamination acting after censoring does not correspond to a meaningful perturbation of the latent distribution, and therefore provides limited insight for robustness assessment in reconstruction-based approaches.
Overall, this contribution opens new perspectives for the development of robust inference tools in censored extreme value models. Promising directions for future research include adaptive threshold selection, extensions to dependent or covariate-dependent censoring mechanisms, and the design of robust estimators capable of handling severe censoring scenarios where the proportion of uncensored extremes is small $(p\leq 1/2).$
Declarations
  • The authors have no findings to declare.
  • The authors have no relevant financial or non-financial interests to disclose.
  • The authors have no competing interests, that are relevant to the content of this article, to declare.
  • All authors certify that they have no affiliations with or involvement in any organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.
  • The authors have no financial or proprietary interests in any material discussed in this article.
7. Acknowledgements
The authors would like to thank the anonymous reviewers for their careful reading of the manuscript and their constructive comments, which helped improve the clarity and quality of this work. We also acknowledge helpful discussions with colleagues that contributed to the development of the theoretical and numerical aspects of the paper. Any remaining errors are the sole responsibility of the authors.